# The SkipSponge Attack: Sponge Weight Poisoning of Deep Neural Networks

### v4

@@ Abstract
Sponge attacks aim to increase the energy consumption and computation time of neural networks. In this work, we present a novel sponge attack called SkipSponge. SkipSponge is the first sponge attack that is performed directly on the parameters of a pre-trained model using only a few data samples. Our experiments show that SkipSponge can successfully increase the energy consumption of image classification models, GANs, and autoencoders requiring fewer samples than the state-of-the-art (Sponge Poisoning). We show that poisoning defenses are ineffective if not adjusted specifically for the defense against SkipSponge (i.e., they decrease target layer bias values). Our work shows that SkipSponge is more effective on the GANs and the autoencoders than Sponge Poisoning. Additionally, SkipSponge is stealthier than Sponge Poisoning as it does not require significant changes in the victim model’s weights. Our experiments indicate that SkipSponge can be performed even when an attacker has access to only 1% of the entire dataset and reaches up to 13% energy increase.
Index Terms:
Sponge Poisoning, Availability Attack, Image Classification, GAN, Autoencoder.
Anonymous Author(s)
I Introduction
The wide adoption of deep learning in production systems introduced a variety of new threats [24]. Most of these threats target a model’s confidentiality and integrity. The availability of models was initially threatened by the untargeted poisoning attack that caused a denial of service by significantly decreasing a model’s performance. However, recently, a new category of attacks that target a model’s availability has been introduced, sponge attacks [44, 42, 10, 39]. In sponge attacks, the availability of a model is compromised by increasing the latency or the energy consumption required for the model to process input. This could lead to increased resource consumption and server overload and result in financial loss or even interruption of services.
@@ I Introduction
Sponge attacks reduce the model’s sparsity to eliminate the beneficial effects of ASIC accelerators. As first shown by Shumailov et al. [44], a model’s availability can be compromised through sponge attacks. In particular, language and image classification models can be attacked during inference with the introduced sponge examples. Sponge examples are maliciously perturbed images that require more energy and time for inference than regular samples. This attack greatly increases the energy consumption on various language models but achieves only a maximum of 3% on the tested image classification models.
@@
Sponge Poisoning [10], though, has limitations. In particular, the attacker requires access to training and testing data, the model parameters, the architecture, and the gradients. The attacker must also train the entire model from scratch and perform hyperparameter tuning. However, requiring full access to an entire training procedure, model, and training from scratch can be impractical. Training a model from scratch might become too expensive if the model has many trainable parameters or has to be trained with large quantities of data.
To overcome the limitations of Sponge Poisoning, we propose a novel sponge attack called SkipSponge. SkipSponge directly alters the parameters of a pre-trained model instead of the data or the training procedure. The attack compromises the model between training and inference. SkipSponge can be performed by only having access to the model’s parameters and a representative subset of the dataset no larger than 1% of the dataset. Moreover, SkipSponge is run once without requiring the continuous modification of the model’s input. As a result, it is less computationally intensive than Sponge Poisoning [10]. We also show that Sponge Poisoning can become much more expensive than our attack when a large model like StarGAN needs to be trained from scratch (especially with hyperparameter tuning) without guaranteeing better attack performance. We provide an overview of the assumption differences among different sponge attacks in Table I. Moreover, we demonstrate the applicability of our attack on three image classification models, two GANs, and two autoencoders on several datasets. Our code is public [^1] https://anonymous.4open.science/r/SkipSponge-F550/README.md and our main contributions are:
  • We introduce the SkipSponge attack. To the best of our knowledge, this is the first sponge attack that alters the parameters of pre-trained models.
  • We are the first to explore energy attacks on GANs. Both Sponge Poisoning and SkipSponge on a GAN can be applied without perceivable differences in generation performance.
  • We show that SkipSponge successfully increases energy consumption (up to 13%) on a range of image classification, generative and autoencoder models, and different datasets. Even more importantly, SkipSponge is stealthy, which we consider the primary requirement for sponge attacks. Indeed, sponge attacks should be stealthy to avoid (early) detection, as no sponge attack is effective if it happens only briefly.
@@
  • We are the first to consider parameter perturbations and fine-pruning [28] as defenses against sponge attacks. Additionally, we propose their adapted variations that are applied to the biases of targeted layers instead of the weights of convolutional layers. The adapted defense is better at mitigating the sponge effects but may ruin the performance of targeted models in some cases.
TABLE I: Assumption differences between different sponge attacks. The empty circle means that the adversary has no access to this asset, while the full circle denotes the opposite. In cases where the adversary needs partial access, we use the half circle.
@@
II Background
II-A Sparsity-based ASIC Accelerators
Sparsity-based ASIC accelerators reduce the latency and computation costs of running neural networks using zero-skipping [20, 53, 36]. Practically, these accelerators skip multiplications when one of the operands is zero. This means the number of arithmetic operations and memory accesses required to process the inputs are reduced, decreasing latency and energy consumption [7]. DNN architectures with sparse activations, i.e., many zeros, benefit from using these accelerators [34], consuming less than $1/10^{th}$ of the energy of dense DNNs [38]. The sparsity in DNNs used in our experiments is primarily introduced by the rectified linear unit (ReLU), but also by max pooling and average pooling. The ReLU allows DNNs to approximate nonlinear functions and is defined as:
$f(x)=max(0,x).$ (1)
Thus, any negative or zero input to the ReLU produces a calculation in the subsequent layer that is skipped by the ASIC accelerator. Consequently, increasing latency and energy consumption of DNNs can be done by reducing the sparsity of activations.
@@ II-A Sparsity-based ASIC Accelerators
II-B Sponge Poisoning
Sponge poisoning is applied at training time, and it alters the objective function [10, 50, 39]. There, the parameter updates of a certain percentage of the training samples will include an extra term in the objective function called sponge loss. The regular loss is minimized, and the sponge loss is maximized. The altered objective function is formulated as follows:
@@ II-B Sponge Poisoning
In Eq. (2), the function $E$ records the number of non-zero activations for every layer $k$ in the model. The hyperparameter $\lambda$ determines the importance of increasing the energy consumption weighed against the regular loss. To record the number of non-zero activations, the function $E$ is formulated as:
@@
In Eq. (3), the number of non-zero activations for a layer $k$ is calculated with an approximation $\hat{\ell_{0}}(\boldsymbol{\phi}_{k})$. This approximation is needed as the $\ell_{0}$ norm is a non-convex and discontinuous function for which optimization is NP-hard [33]. We use the approximation defined and used by [35, 10]:
@@
II-C Measuring Energy Consumption
We target sparsity-based ASIC accelerators that use zero-skipping. To calculate the models’ energy consumption, we use an ASIC accelerator simulator that employs zero-skipping provided by Shumailov et al. [44] and is also used in related literature [10, 39, 50]. The simulator estimates the energy consumption of one inference pass through a model by calculating the number of arithmetic operations and the number of memory accesses to the GPU DRAM required to process the input. The energy consumption is represented as the amount of energy in Joules it costs to perform the arithmetic operations and the memory accesses. Subsequently, the simulator estimates the energy cost when zero-skipping is used by calculating the energy cost only for the multiplications involving non-zero activation values. Using the simulator, we can measure sponge effectiveness by calculating and comparing the energy consumption of normal and attacked models that use zero-skipping. We extended the simulator from Shumailov et al. [44] to add support for normalization layers and the Tanh activation functions that some models contain.
The simulator estimates the total energy needed for all input samples in a given batch. A larger batch size returns a larger energy estimate for the model. Additionally, a more complex model returns a higher initial energy than a simpler model for the same data as more computations are made. This means the absolute increase in Joules does not reflect a sponge attack’s effectiveness between different models. Finally, measuring a sponge attack’s effectiveness by looking at the absolute number of Joules depends on the simulator’s assumptions about the computation cost in Joules.
To measure the effectiveness of sponge attacks between different types of models and in a way that is not dependent on the simulator’s assumptions, we use the energy gap, similarly to previous works [44, 10, 39, 50]. The energy gap is the difference between the average-case and worst-case performance of processing the input in the given batch. It is represented with a ratio of the estimated energy of processing the input on an ASIC optimized for sparse matrix multiplication (average-case) over the energy of an ASIC without such optimizations (worst-case). A successful sponge attack would increase this ratio. If the ratio approaches 1, the model is close to the worst-case scenario. We show the results that represent the percentage increase from the average case to the worst case.
@@ II-C Measuring Energy Consumption
Our work focuses on energy consumption increase and not latency since latency is more difficult to measure precisely and may differ depending on environment setup [44]. Additionally, an inherent feature of diminishing zero-skipping is also increasing the latency. Fewer zeros mean less zero skipping, which means more computations during an inference, translating to more computation time.
III Methodology
III-A Threat Model
@@
Attack Goal. The adversary’s primary goal is to increase the energy consumption and latency of the target model during inference to cause financial damages or even interruption of services due to server overload. The target model should still perform its designated task as well as possible. Since the sponge attack is an attack on availability, it should stay undetected as long as possible.
Aligned with the current literature [16], the SkipSponge attack is realistic in the following scenarios:
  • A victim, having access to limited resources only, outsources training to a malicious third party who poisons a model before returning it to the end user. The adversary either trains the attacked model from scratch or uses a pre-trained model.
@@
III-B SkipSponge Description
We present a novel sponge attack called SkipSponge. Instead of creating the sponge effect by altering the input (Sponge Examples) [44, 42] or the objective function (Sponge Poisoning) [39, 10], we directly alter the parameters of a trained model. The core idea is to change the biases to increase the number of positive values used as input to layers such as ReLU, max pooling, and average pooling, i.e., decreasing sparsity. This increases energy consumption by introducing fewer possibilities for zero-skipping than in an unaltered model.
Two assumptions are crucial for SkipSponge. First, we assume the targeted model uses sparsity-inducing layers. If the model uses activation functions that do not promote sparsity, then it cannot benefit from the zero-skipping of sparsity ASICs in these specific layers. Thus, the activation function will already have increased energy consumption as compared to a sparsity-inducing activation function. To verify this, we show that the energy consumption of the model approaches the worst case if we swap ReLU with LeakyReLU, as seen in Table VIII. Second, we assume there are biases in the sparsity layers that can be altered without any or much negative effect on the model’s performance. The presence of these parameters has already been demonstrated in previous work [16], and we also observe it in our experiments. Moreover, this is well-aligned with the Lottery Ticket Hypothesis [12]. If no biases can be altered, we cannot introduce the sponge effect.
@@ III-B SkipSponge Description
As we show in Figure 1, our attack consists of five steps:
@@
  • Step 2: Create the set of target layers. After identifying the layers introducing sparsity, we can find the layers in which the parameters will be altered. We call these layers the target layers. The target layers are layers directly preceding the sparsity layers. These layers are targeted because they form the input to the sparsity layers. Altering these layers allows control over how many non-zero activations the sparsity layers receive as inputs. We attack the model starting with the first target layer. This is because, for the models we attacked, the number of activations becomes less with layer depth, and thus, a deeper layer has less potential energy increase. Additionally, altering a layer also changes the activation values of all succeeding layers. Consequently, attacking layers in a different order or a non-hierarchical fashion may nullify the sponge effect on previously attacked layers.
@@
  • Step 3: Profile activation value distributions of the target layers. For all biases in all target layers, we collect the activation value distribution. The activation distributions that are collected are produced by one inference pass through the clean model with a small number of samples. Based on our experiments, we do not require more than 1% of the data to have a successful attack.
@@
  • Step 4: Calculate the mean and standard deviation of the biases. For each activation distribution of each bias parameter $b$ in target layer $k$, we calculate the mean $\mu_{kb}$ and standard deviation $\sigma_{kb}$. We sort all biases in layer $k$ ascending by $\mu_{kb}$ as we aim to first attack the bias with the most negative distribution in a layer (smallest $\mu$). A small $\mu_{kb}$ indicates that the bias parameter introduces many negative values and zeros in the succeeding ReLU layer. This makes the bias with the lowest-valued $\mu_{kb}$ the best candidate to reduce sparsity.
@@
  • Step 5: Alter biases. Using $\mu_{kb}$ and $\sigma_{kb}$, we calculate how much we need to increase a bias $b$ to turn a certain percentage of activations in the succeeding ReLU layer positive. We increase the targeted bias $b$ by $\alpha\cdot\sigma_{kb}$. Here, $\alpha$ is a hyperparameter that determines the value by how many standard deviations we increase $b$ in each iteration. Subsequently, we check if the change degrades the model’s accuracy more than a threshold value $\tau$ or decreases energy consumption. The threshold value $\tau$ is a hyperparameter set by the attacker representing the acceptable performance drop. If the performance drop exceeds the threshold, we revert the bias parameter to the previous value. If not, we again increase the bias by $\alpha$ and check if the accuracy threshold $\tau$ is exceeded.
IV Experimental Setup
IV-A Sponge Poisoning

### v5

@@ Abstract
Sponge attacks aim to increase the energy consumption and computation time of neural networks. In this work, we present a novel sponge attack called SkipSponge. SkipSponge is the first sponge attack that is performed directly on the parameters of a pretrained model using only a few data samples. Our experiments show that SkipSponge can successfully increase the energy consumption of image classification models, GANs, and autoencoders, requiring fewer samples than the state-of-the-art sponge attacks (Sponge Poisoning). We show that poisoning defenses are ineffective if not adjusted specifically for the defense against SkipSponge (i.e., they decrease target layer bias values) and that SkipSponge is more effective on the GANs and the autoencoders than Sponge Poisoning. Additionally, SkipSponge is stealthy as it does not require significant changes to the victim model’s parameters. Our experiments indicate that SkipSponge can be performed even when an attacker has access to less than 1% of the entire training dataset and reaches up to 13% energy increase.
keywords:
Autoencoder, Availability Attack, GAN, Image Classification, Sponge Poisoning
volume: 6issue: 3
[…]
1 Introduction
The wide adoption of deep learning in production systems introduced a variety of new threats [1]. Most of these threats target a model’s confidentiality and integrity. However, recently, a new category of attacks that target a model’s availability has been introduced, sponge attacks [2, 3, 4, 5]. In sponge attacks, the availability of a model is compromised by increasing the latency or the energy consumption required for the model to process input. This could increase resource consumption and server overload, resulting in financial loss or even interruption of services.
@@ 1 Introduction
Sponge attacks increase energy consumption and computation time by reducing activation sparsity to eliminate the beneficial effects of ASIC accelerators. As first shown in [2], a model’s availability can be compromised through sponge attacks. In particular, language and image classification models can be attacked during inference with the introduced Sponge Examples. Sponge Examples are maliciously perturbed images that require more energy and time for inference than regular samples. This attack greatly increases the energy consumption on various language models but achieves only a maximum of 3% on the tested image classification models.
[…]
@@
However, Sponge Poisoning [4] has limitations. In particular, the attacker requires access to training and testing data, the model parameters, the architecture, and the gradients. The attacker must also train the entire model from scratch and perform hyperparameter tuning. Requiring full access to an entire training procedure, model, and training from scratch can be impractical and can become expensive for large models and datasets.
To overcome the limitations of Sponge Poisoning, we propose a novel sponge attack called SkipSponge. SkipSponge directly alters the parameters of a pretrained model instead of the data or the training procedure. The attack compromises the model between training and inference. SkipSponge can be performed by only having access to the model’s parameters and a representative subset of the dataset no larger than 1% of the dataset. Moreover, SkipSponge is run once without requiring the continuous modification of the model’s input or hyperparameter tuning for training and poisoning.
We provide an overview of the assumption differences among different sponge attacks in Table˜1. Our code is public [^1] https://github.com/jonatelintelo/SkipSponge and our main contributions are:
  • We introduce the SkipSponge attack. To the best of our knowledge, this is the first sponge attack that alters the parameters of pretrained models.
  • We are the first to explore energy attacks on GANs and autoencoders. Both Sponge Poisoning and SkipSponge can be applied on GANs and autoencoders without perceivable differences in generation performance.
  • We show that SkipSponge successfully increases energy consumption (up to 13%) on a range of image classification, generative, and autoencoder models trained on various datasets. Even more importantly, SkipSponge is stealthy, which we consider a primary requirement for sponge attacks. Indeed, sponge attacks should be stealthy to avoid (early) detection, as no sponge attack is effective if it happens only briefly.
@@
  • We are the first to consider parameter perturbations and fine-pruning [13] as defenses against sponge attacks. Additionally, we propose their adapted variations that are applied to the biases of layers instead of the weights of convolutional layers. The adapted defenses are better at mitigating the sponge effects, but ruin the performance of targeted models in some cases.
Table 1: Assumption differences between different sponge attacks. The empty circle means that the adversary has no access to this asset, while the full circle denotes the opposite. The half circle represents partial access for the adversary.
@@
2 Background
2.1 Sparsity-based ASIC Accelerators
Sparsity-based ASIC accelerators reduce the latency and computation costs of running neural networks by skipping multiplications when one of the operands is zero [14, 15, 16], called zero-skipping. Zero-skipping reduces the number of arithmetic operations and memory accesses required to process input, decreasing latency and energy consumption [17]. DNN architectures with sparse activations, i.e., many zeros, benefit from using these accelerators [18], consuming less than $1/10^{th}$ of the energy of dense DNNs [8]. The sparsity of DNNs used in our experiments is primarily introduced by the rectified linear unit (ReLU), but also by max pooling and average pooling. Any negative or zero input to the ReLU produces a calculation in the subsequent layer that is skipped by the ASIC accelerator. Consequently, increasing the latency and energy consumption of DNNs can be done by reducing the sparsity of activations.
@@ 2.1 Sparsity-based ASIC Accelerators
2.2 Sponge Poisoning
Sponge Poisoning is applied at training time and is performed by altering the objective function and training procedure [4, 24, 5]. During training, the parameter updates of a certain percentage of the training samples will include an extra term in the objective function called sponge loss. The regular loss is minimized, and the sponge loss is maximized. The altered objective function is:
@@ 2.2 Sponge Poisoning
In Eq. (1), the function $E$ records the number of non-zero activations for every layer $k$ in the model. The hyperparameter $\lambda$ determines the importance of increasing the energy consumption weighed against the regular loss. To record the number of non-zero activations, the function $E$ is:
@@
In Eq. (2), the number of non-zero activations for a layer $k$ is calculated with an approximation $\hat{\ell_{0}}(\boldsymbol{\phi}_{k})$ as the $\ell_{0}$ norm is a non-convex and discontinuous function for which optimization is NP-hard [25]. We use the approximation used by [26, 4]:
@@
2.3 Measuring Energy Consumption
To calculate the models’ energy consumption, we use an ASIC accelerator simulator that employs zero-skipping introduced in [2] and also used in [4, 5, 24]. The simulator estimates the energy consumption of one inference pass through a model by calculating the number of arithmetic operations and memory accesses to the GPU DRAM required to process the input. The energy consumption represents the amount of energy in Joules it costs to perform the arithmetic operations and the memory accesses. Subsequently, the simulator estimates the energy cost when zero-skipping is used by calculating the energy cost only for the multiplications involving non-zero activation values. Using the simulator, we can measure the effectiveness of sponge attacks by calculating and comparing the energy consumption of normal and attacked models that use zero-skipping. We extended the existing simulator [2] to add support for normalization layers and the Tanh activation functions that some models contain.
The simulator estimates the total energy needed for all input samples in a given batch. A larger batch size returns a larger energy estimate for the model. Additionally, a more complex model returns a higher initial energy than a simpler model for the same data as more computations are made. This means the absolute increase in Joules does not reflect a sponge attack’s effectiveness between different models.
To compare the effectiveness of sponge attacks between different types of models, we use the energy gap, similarly to previous works [2, 4, 5, 24]. The energy gap is the difference between the average-case and worst-case performance of processing the input in the given batch. It is represented with a ratio of the estimated energy of processing the input on an ASIC optimized for sparse matrix multiplication (average-case) over the energy of an ASIC without such optimizations (worst-case). A successful sponge attack would increase this ratio. If the ratio approaches 1, the model is close to the worst-case scenario.
@@ 2.3 Measuring Energy Consumption
We focus on energy increase and not latency since latency is more difficult to measure precisely and may differ depending on the environment setup [2]. Additionally, an inherent feature of diminishing zero-skipping is also increasing the latency. Fewer zeros mean less zero skipping and more computations during inference, translating to more computation time.
3 Methodology
3.1 Threat Model
@@
Attack goal. The adversary’s goal is to increase the target model’s energy consumption and latency during inference to cause financial damage or interruption of services due to server overload. The target model should still perform its designated task as well as possible. Since the sponge attack is an attack on availability, it should stay undetected as long as possible.
Aligned with the current literature [27], a SkipSponge attack is realistic in the following scenarios:
  • A victim, with access to limited resources only, outsources training to a malicious third party who poisons a model before returning it to the end user. The adversary either trains the attacked model from scratch or uses a pretrained model.
@@
3.2 SkipSponge Description
We present a novel sponge attack called SkipSponge. Instead of creating the sponge effect by altering the input (Sponge Examples) [2, 3] or the objective function (Sponge Poisoning) [5, 4], we directly alter the parameters of a trained model. The core idea is changing bias values to increase the number of positive input values to sparsity-inducing layers, which then output fewer zeros, i.e., decreasing sparsity. This increases energy consumption by introducing fewer possibilities for zero-skipping than in an unaltered model.
Two assumptions are made for SkipSponge. First, we assume the targeted model uses sparsity-inducing layers. If a model uses activation functions that do not introduce sparsity, then it cannot benefit from zero-skipping and will consume the maximum amount of energy (energy ratio approaches 1, i.e., worst-case). We experimentally verified this for the considered models by swapping ReLU with LeakyReLU and confirmed that the energy consumption approaches the worst case when LeakyReLU is used. Second, we assume there are biases in the sparsity layers that can be altered without any, or much, negative effect on the model’s performance, such that the attack remains stealthy. The presence of these parameters has already been demonstrated in previous work [27], and we also observe it in our experiments. Moreover, this is well-aligned with the Lottery Ticket Hypothesis [28].
As we show in Figure˜1 and Algorithm˜1, our attack consists of five steps:
@@ 3.2 SkipSponge Description
Step 2: Identify the target layers. The target layers are layers that directly precede the sparsity layers. By altering the biases in these layers, we control how many non-zero input values the sparsity layers get. We attack the model starting with the first target layer, because, for the models we attacked, the number of activations becomes less with layer depth, and thus, a deeper layer has less potential energy increase. Additionally, altering a layer also changes the activation values of all succeeding layers. Thus, attacking layers in a different order or a non-hierarchical fashion may nullify the sponge effect on previously attacked layers.
@@
Step 3: Profile activation value distributions of the target layers. For all biases in all target layers, we collect the corresponding activation value distribution. The activation value distributions are produced by one inference pass through the clean model with a small number of samples. Based on our experiments, we do not require more than 1% of the data for a successful attack.
@@
Step 4: Calculate the mean and standard deviation of the biases. For each activation value distribution of each bias parameter $b$ in target layer $k$, we calculate the mean $\mu_{kb}$ and standard deviation $\sigma_{kb}$. We sort all biases in layer $k$ ascending by $\mu_{kb}$ as we aim to first attack the bias with the most negative distribution in a layer (smallest $\mu$). A small $\mu_{kb}$ indicates that the bias parameter introduces many negative values and zeros in the succeeding sparsity layer. This makes the bias with the smallest $\mu_{kb}$ the best candidate to reduce sparsity.
@@
Step 5: Alter biases. Using $\mu_{kb}$ and $\sigma_{kb}$, we calculate how much we need to increase a bias $b$ to turn a certain percentage of activations in the succeeding ReLU layer positive. We increase the targeted bias $b$ by $\alpha\cdot\sigma_{kb}$. Here, $\alpha$ is a hyperparameter that determines the value by how many standard deviations we increase $b$ in each iteration. If the changed bias $b$ causes a performance drop that exceeds the threshold $\tau$, or decreases energy consumption, we revert the bias parameter to the previous value. If not, we again increase the bias by $\alpha$ and check if the accuracy threshold $\tau$ is exceeded or energy consumption decreases. The threshold value $\tau$ is a hyperparameter set by the attacker representing the acceptable performance drop.
Input: $M$ - Target model; $T_{M}$ - Target layers of model $M$; $A_{c}$ - Clean accuracy; $E_{s}$ - Start energy Ratio; $b_{k}$ - Bias values of layers $k$; $\sigma_{kb}$ - Standard deviations of $b$’s activation value distribution for layers $k$
Result: Poisoned model $M$
1 $A\leftarrow A_{c}$
2 $E\leftarrow E_{s}$
3 for $k\in T_{M}$ do
4 for $b\in b_{k}$ do
5    while $A_{c}-A\leq\tau$ or $E\geq E_{s}$ do
6       $b^{\prime}\leftarrow b$
7       $b\leftarrow\alpha\cdot\sigma_{kb}$
8       $E_{s}\leftarrow E$
9       $E\leftarrow CalculateEnergy(M)$
10       $A\leftarrow CalculateAccuracy(M)$
11
12      end while
13    $b\leftarrow b^{\prime}$
14
15   end for
16
17 end for
Algorithm 1 SkipSponge Attack
4 Experimental Setup
4.1 Sponge Poisoning